# Thera: Aliasing-Free Arbitrary-Scale Super-Resolution with Neural Heat Fields

### v3

@@ 3.3 Total Variation at $\bm{t=0}$
To allow Thera to better generalize to higher, out-of-domain scaling factors, we can place an unsupervised regularizer at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well [11]. We use an $\ell^{1}$ variant of TV,
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work [38]. We further motivate this approach in Fig. 5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
@@ 3.4 Implementation and Training
Thera is implemented in JAX [6]. Similar to prior work [10, 24, 7, 56], we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer [23] with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The learning rate is decayed to zero according to a cosine annealing schedule [30]. We use MAE as reconstruction loss, to which the TV loss from Eq. 5 is added with a weight of $10^{-4}$. Like previous work [43, 26, 44], we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF [10]. In GSE, the results for four rotated versions of the input are averaged at test time. Including reflections did not improve performance.
@@ 4 Results
  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based [29] hypernetwork. Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like [7].
  • Thera Pro: The strongest version uses a high-capacity, transformer-based [28, 25] hypernetwork. Its added parameter count is $\approx$4.63 M, still less than the most recent competitor [56] and much smaller than [9].
Datasets and metrics. Following previous work, our models are trained with the DIV2K [1] training set, consisting of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 [5], Set14 [53], BSDS100 [31], Urban100 [19], and Manga109 [32]. Following prior work, we use peak signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric and compute it in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Additional quantitative results are given in Appendix C. Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix F).
Backbones. Like previous work, we combine our method with two standard backbones for super-resolution and image restoration: (i) EDSR-baseline [26] (1.22 M parameters) and (ii) RDN [54] (22.0 M parameters).
@@ 4.1 Super-Resolution Performance
Low-res input LIIF [10] SRNO [45] MSIT [56] Thera Pro (ours) GT
DIV2K
DIV2K
Set14
Urban100
Manga109
Manga109
Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN [54] backbone for all methods. Best viewed zoomed in.
Qualitative results. Upon visual inspection – see Fig. 4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures. Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Fig. 9 in Appendix B).
@@
EDSR- baseline [26] (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
MetaSR [18] 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
LIIF [10] 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
LTE [24] 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
CUF [44] 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
CiaoSR [7] 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
CLIT [9] 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
SRNO [45] 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
MSIT [56] 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
@@
RDN [54] (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
MetaSR [18] 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
LIIF [10] 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
LTE [24] 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
CUF [44] 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
CiaoSR [7] 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
CLIT [9] 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
SRNO [45] 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
MSIT [56] 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
Thera Air .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
Thera Plus 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
Thera Pro 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
Quantitative results. We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above. Table 1 shows PSNR values for all tested methods, for both in-distribution (×$\times 2$ to ×$\times 4$) and out-of-distribution (×$\times 6$ to ×$\times 30$) scaling factors. Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs. 29.22 on EDSR ×$\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT [56], and less than a third of CLIT [9]. Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count. This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction. For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective ×$\times 4$ backbone. This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
@@
Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN [54] backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Tab. 2. Our method again outperforms all competing methods in all settings, often substantially (e.g., 29.58 vs. 29.14 on Urban100 ×$\times 3$). We hypothesize that Thera’s hard-wired PSF is also beneficial when generalizing to unseen datasets. Once again we observe that the performance of Thera Air is often comparable to that of methods with orders of magnitude higher parameter counts.
Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor [10]. Due to its natural, Fourier-inspired representation, Thera is also infinitely differentiable, while ReLU-based competitors approximate the signal as a piecewise-linear function with null higher derivatives (last row).
Fidelity of the Signal and its Derivatives. Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives [41]. We observe similar effects for Thera, whose thermal activations at $t=0$ can be seen as a special case of periodic activations, cf. Figure 5. In fact, due to the use of thermal activations – and unlike all prior work based on multi-layer ReLU-activated fields – Thera is infinitely differentiable.
@@ 4.2 Ablation Studies
$\times$2 only, fixed $\kappa$ 35.01 31.00 28.84 26.47 21.97 20.33
$\times$3 only, fixed $\kappa$ 34.74 31.25 29.25 26.91 22.25 20.54
$\times$4 only, fixed $\kappa$ 34.47 31.17 29.24 26.97 22.32 20.59
Fixed $\kappa$ 34.86 31.22 29.24 26.95 22.31 20.58
@@
Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
@@
Single scale training. We run three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization. $\kappa$ was fixed at the theoretically derived value for these experiments, as multi-scale training is required to optimize it. As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Tab. 3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
Trainable $\bm{\kappa}$. Fixing $\kappa$ at the theoretically derived value (Eq. 3) leads to a small drop in performance. This suggests that there remain effects that are not accounted for by our proposed observation model, albeit very minor.
Geometric self-ensemble. In line with previous work [43, 26, 44] we see a notable performance boost with geometric self-ensembling. Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
@@
Thermal activations. We replace thermal activations (Eq. 2) with standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
@@ A.1 Preliminaries
In this formulation, we attempt to recover a “pure” signal at $t=0$ or higher sampling rates $10<t<$ 1 given an observation at $t=1$. Note that
$\Phi(\mathbf{x},t)\text{ is}\begin{cases}\text{meaningless},&\text{if}\ t<0\\ \text{pure signal},&\text{if}\ t=0\\ \text{ill-posed},&\text{if}\ 0<t<1\\ I,&\text{if}\ t=1\\ \text{well-posed},&\text{if}\ t>1\\ \end{cases}$ (9)
The ill-posed problem for $10<t<$ 1 is the interesting case, where this formulation relates to super-resolution. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
@@ A.2 Thermal Diffusivity Coefficient
To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$. One way to do so is to match the cut-off frequency of the filter in Eq. 7 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate. We take the cut-off frequency of the Gaussian filter defined in Eq. 7 to be the frequency whose amplitude is halved, which is
@@
For the signal compressed into the domain $[-0.5,0.5]$, we can compute the Nyquist frequency to be
@@
Finally, we would like to highlight that Eq. 10 is specific to the case where $\mathbf{x}$ is 2-dimensional. The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using our field’s formulation in spaces with numbers of spatial dimensions other than 2. Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the formulas for a Gaussian filter with the appropriate number of dimensions.
@@ A.3 Relationship Between t and s
Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate. If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it down to $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$. Using Eq. 10 and Eq. 11, we can easily find the quadratic relationship
@@
For instance, if we want to upsample the image by a factor of 2, we should use $t=0.5^{2}=0.25$. Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling. This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
@@
Figure 6 shows an example where we fit a neural heat field at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 4. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
@@
Figure 8: Blurring an image before a point-wise sample is equivalent to observing with PSF equivalent to the blur kernel.
In Fig. 7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations. Note that aliasing is not always as obvious to the eye as Moiré patterns: in the cases shown in the figure, it would simply mean that sampling the center location at $t=0$ would not be representative of the pixel’s footprint. Figure 8 further illustrates how such blur is equivalent to a scale-appropriate anti-aliasing filter.

### v4

@@ 3.3 Total Variation at $\bm{t=0}$
To allow Thera to better generalize to higher, out-of-domain scaling factors, we can place an unsupervised regularizer at $t=0$. Note that this is a prior on the continuous signal itself – something that, to the best of our knowledge, sets Thera apart from all previous methods. In our implementation it takes the form of a total variation (TV) loss term, well known to promote piece-wise constant signals that describe natural images well (Chugunov et al., 2024). We use an $\ell^{1}$ variant of TV,
@@
Given our continuous signal representation, $\nabla\Phi(\mathbf{x},0)$ can be computed analytically by automatic differentiation, rather than falling back to a neighborhood approximation as in most previous work (Rudin et al., 1992). We further motivate this approach in Figure 5, which demonstrates that our method faithfully recovers the gradients of super-resolved images.
@@ 3.4 Implementation and Training
Thera is implemented in JAX (Bradbury et al., 2018). Similar to prior work (Chen et al., 2021; Lee & Jin, 2022; Cao et al., 2023; Zhu et al., 2025), we randomly sample a scaling factor $r\sim\mathrm{\mathcal{U}(1.2,4)}$ for each image during training, then randomly crop an area of size $(48r)^{2}$ pixels as the target patch, from which the source is generated by bicubic downsampling to size $48^{2}$. As corresponding targets, $48^{2}$ random pixels are sampled from the target patch. We train with standard augmentations (random flipping, rotation, and resizing), using the Adam optimizer (Kingma & Ba, 2015) with a batch size of $16$ for $5\times 10^{6}$ iterations, with initial learning rate $10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The learning rate is decayed to zero according to a cosine annealing schedule (Loshchilov & Hutter, 2016). We use MAE as reconstruction loss, to which the TV loss from Eq. 5 is added with a weight of $10^{-4}$. Like previous work (Timofte et al., 2016; Lim et al., 2017; Vasconcelos et al., 2023), we employ geometric self-ensembling (GSE) instead of the local self-ensembling introduced in LIIF (Chen et al., 2021). In GSE, the results for four rotated versions of the input are averaged at test time. Including reflections did not improve performance.
@@ 4 Results
  • Thera Plus: A balanced version that employs an efficient ConvNeXt-based (Liu et al., 2022) hypernetwork. Its parameter count of $\approx$1.41 M matches that of recent medium-sized competitors like Cao et al. (2023).
  • Thera Pro: The strongest version uses a high-capacity, attention-based hypernetwork. Its added parameter count is $\approx$4.63 M, still less than the most recent competitor (Zhu et al., 2025) and much smaller than Chen et al. (2023), both attention-based.
Datasets and metrics. Following previous work, our models are trained with the DIV2K (Agustsson & Timofte, 2017) training set, consisting of 800 high-resolution RGB images of diverse scenes. We report evaluation metrics on the official DIV2K validation split as well as on standard benchmark datasets: Set5 (Bevilacqua et al., 2012), Set14 (Zeyde et al., 2012), BSDS100 (Martin et al., 2001), Urban100 (Huang et al., 2015), and Manga109 (Matsui et al., 2017). Following prior work, we use peak signal-to-noise ratio (PSNR, in decibels) as the main evaluation metric and compute it in RGB space for DIV2K and on the luminance (Y) channel of the YCbCr representation for benchmark datasets. Additional quantitative results are given in Appendix C. Not all numbers could be computed for competing methods for which code or checkpoints were not publicly shared (see Appendix H).
Backbones. We combine each of the three variants of our method with two standard backbones for super-resolution and image restoration, as done in previous work: (i) EDSR-baseline (Lim et al., 2017) (1.22 M parameters) and (ii) RDN (Zhang et al., 2018) (22.0 M parameters).
@@ 4.1 Super-Resolution Performance
EDSR- baseline (1.22 M) + Sinc interpol. 34.52 30.89 28.98 26.75 23.76 22.25 21.29 20.62
MetaSR 0.45 M 34.64 30.93 28.92 26.61 23.55 22.03 21.06 20.37
LIIF 0.35 M 34.67 30.96 29.00 26.75 23.71 22.17 21.18 20.48
LTE 0.49 M 34.72 31.02 29.04 26.81 23.78 22.23 21.24 20.53
CUF 0.30 M 34.79 31.07 29.09 26.82 23.78 22.24
CiaoSR 1.43 M 34.88 31.12 29.19 26.92 23.85 22.30 21.29 20.44
CLIT 15.7 M 34.81 31.12 29.15 26.92 23.83 22.29 21.26 20.53
SRNO 0.80 M 34.85 31.11 29.16 26.90 23.84 22.29 21.27 20.56
MSIT 4.83 M 34.95 31.23 29.22 26.94 23.83 22.27 21.26 20.54
@@
RDN (22.0 M) + Sinc interpol. 34.59 31.03 29.12 26.89 23.87 22.34 21.36 20.68
MetaSR 0.45 M 35.00 31.27 29.25 26.88 23.73 22.18 21.17 20.47
LIIF 0.35 M 34.99 31.26 29.27 26.99 23.89 22.34 21.31 20.59
LTE 0.49 M 35.04 31.32 29.33 27.04 23.95 22.40 21.36 20.64
CUF 0.30 M 35.11 31.39 29.39 27.09 23.99 22.42
CiaoSR 1.43 M 35.13 31.39 29.43 27.13 24.03 22.45 21.41 20.55
CLIT 15.7 M 35.10 31.39 29.39 27.12 24.01 22.45 31.38 20.64
SRNO 0.80 M 35.16 31.42 29.42 27.12 24.03 22.46 21.41 20.68
MSIT 4.83 M 35.16 31.42 29.42 27.11 23.99 22.42 21.37 20.65
Thera Air (ours) .008 M 35.06 31.42 29.43 27.13 24.04 22.48 21.44 20.71
Thera Plus (ours) 1.41 M 35.00 31.40 29.44 27.16 24.06 22.49 21.45 20.71
Thera Pro (ours) 4.63 M 35.25 31.56 29.57 27.25 24.14 22.56 21.52 20.77
Table 2: Results on common benchmark datasets for in-distribution scale factors with an RDN (Zhang et al., 2018) backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
@@
Quantitative results. We first evaluate the three variants of our method on the held-out DIV2K validation set, following the setup described above. Table 1 shows PSNR values for all tested methods, for both in-distribution ($\times 2$ to $\times 4$) and out-of-distribution ($\times 6$ to $\times 30$) scaling factors. Thera Pro outperforms all competing methods at all scaling factors, often by a substantial margin (e.g., 29.51 vs. 29.22 on EDSR $\times 4$), even though its parameter overhead on top of the backbone is lower compared to the second-best method MSIT (Zhu et al., 2025), and less than a third of CLIT (Chen et al., 2023). Interestingly, even our minimal variant Thera Air – with only about 8000 parameters on top of the backbone – performs on par with or better than methods of much higher parameter count. This supports our claim that hard-wiring a theoretically principled sampling model, which rules out signal aliasing, enables better generalization and higher-fidelity reconstruction. For comparison with conventional interpolation, we also report numbers obtained with Lanczos (sinc) resampling on top of the respective $\times 4$ backbone. This baseline is consistently outperformed by dedicated ASR methods, indicating that the latter do learn scale-specific priors.
Like earlier work, we further report the performance of Thera on five popular benchmark datasets with an RDN backbone in Table 2. Our method again outperforms all competing methods in all settings, often substantially (e.g., 29.58 vs. 29.14 on Urban100 $\times 3$). We hypothesize that Thera’s hard-wired PSF is also beneficial when generalizing to unseen datasets. Once again we observe that the performance of Thera Air is often comparable to that of methods with orders of magnitude higher parameter counts.
Low-res input LIIF SRNO MSIT Thera Pro (ours) GT
DIV2K
DIV2K
Set14
Urban100
Manga109
Manga109
Figure 4: Qualitative examples for a representative $\times$6 scale factor, with an RDN (Zhang et al., 2018) backbone for all methods. Best viewed zoomed in.
Qualitative results. Upon visual inspection – see Figure 4 for examples – we observe that Thera produces results that are both perceptually convincing and more correct, particularly in the presence of repeating structures. Neural heat fields enable Thera to reproduce a high level of detail without suffering from aliasing, no matter the sampling scale (see also Figure 9 in Appendix B).
Figure 5: Thera reconstructs a signal $\Phi$ and its gradient $\nabla_{\mathbf{x}}\Phi$ more faithfully than a ReLU-based competitor (Chen et al., 2021). Due to its natural, Fourier-inspired representation, Thera is also infinitely differentiable, while ReLU-based competitors approximate the signal as a piecewise-linear function with null higher derivatives (last row).
Fidelity of the Signal and its Derivatives. Neural fields with periodic activation functions have been shown to be superior when it comes to fitting high-resolution, natural signals, and to correctly recovering their derivatives (Sitzmann et al., 2020b). We observe similar effects for Thera, whose thermal activations at $t=0$ can be seen as a special case of periodic activations, cf. Figure 5. In fact, due to the use of thermal activations – and unlike all prior work based on multi-layer ReLU-activated fields – Thera is infinitely differentiable.
@@ 4.2 Ablation Studies
Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
@@
$\times$2 only, fixed $\kappa$ 35.00 30.96 28.76 26.34 21.87 20.25
$\times$3 only, fixed $\kappa$ 34.70 31.25 29.25 26.90 22.23 20.53
$\times$4 only, fixed $\kappa$ 34.36 31.18 29.25 26.98 22.31 20.58
Fixed $\kappa$ 34.85 31.22 29.24 26.95 22.29 20.56
@@
Single scale training. We run three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization. $\kappa$ was fixed at the theoretically derived value for these experiments, as multi-scale training is required to optimize it. As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Table 3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
Trainable $\bm{\kappa}$. Fixing $\kappa$ at the theoretically derived value (Equation 3) leads to a small drop in performance. This suggests that there remain effects that are not accounted for by our proposed observation model, albeit very minor.
Geometric self-ensemble. In line with previous work (Timofte et al., 2016; Lim et al., 2017; Vasconcelos et al., 2023) we see a notable performance boost with geometric self-ensembling. Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
@@
Thermal activations. We replace thermal activations (Equation 2) with standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
@@ A.1 Preliminaries
In this formulation, we attempt to recover a “pure” signal at $t=0$ or higher sampling rates $0<t<1$ given an observation at $t=1$. Note that
[…]
The ill-posed problem for $0<t<1$ is the interesting case, where this formulation relates to super-resolution. The super-resolution algorithm should somehow condition the solution space to find the appropriate solution in this domain.
@@ A.2 Thermal Diffusivity Coefficient
To use the above formulations, we need to compute the thermal diffusivity coefficient $\kappa$. One way to do so is to match the cut-off frequency of the filter in Equation 7 at $t=1$ to the well-known Nyquist frequency given by the image’s sampling rate. We take the cut-off frequency of the Gaussian filter defined in Equation 7 to be the frequency whose amplitude is halved, which is
@@
For the signal compressed into the domain [0.5,0.5][-0.5,0.5], we can compute the Nyquist frequency to be
@@
Finally, we would like to highlight that Equation 10 is specific to the case where $\mathbf{x}$ is 2-dimensional. The theoretically ideal value of $\kappa$ is the only part of our formulation that does not directly apply when using our field’s formulation in spaces with numbers of spatial dimensions other than 2. Nonetheless, computing $\kappa$ for other cases would be a simple matter of repeating the steps above with the formulas for a Gaussian filter with the appropriate number of dimensions.
@@ A.3 Relationship Between t and s
Assuming that the field is learned appropriately, we still need to know at what time $t$ we should sample from to obtain the correct (aliasing-free) signal for a different sampling rate. If we define $S$ to be the subsampling rate (i.e., if our base image has $N=128$ and we want to subsample it down to $N=64$, we have $S=2$) we need to find $t$ such that $f_{\text{Nyquist}}$ scales by $1/S$. Using Equation 10 and Equation 11, we can easily find the quadratic relationship
@@
For instance, if we want to upsample the image by a factor of 2, we should use $t=0.5^{2}=0.25$. Thus, $0<t<1$ refers to super-resolution, while $t>1$ refers to downsampling. This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
@@
Figure 6 shows an example where we fit a neural heat field at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Equation 4. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
@@
Figure 8: Blurring an image before a point-wise sample is equivalent to observing with a PSF equivalent to the blur kernel.
In Figure 7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations. Note that aliasing is not always as obvious to the eye as Moiré patterns: in the cases shown in the figure, it would simply mean that sampling the center location at $t=0$ would not be representative of the pixel’s footprint. Figure 8 further illustrates how such blur is equivalent to a scale-appropriate anti-aliasing filter.